OpenStudy (anonymous):

Quadrilateral EFGH is a rectangle. if EF = 4x - 12, HG=2x + 6 , and FG = 2x , find its perimeter.

OpenStudy (anonymous):

let us illustrate it first... 

OpenStudy (anonymous):

since this is a rectangle, all 4 angles are equal, then parallel sides are equal such that EF=HG, as well as EH=FG.... so for EF=HG 4x-12=2x+6 4x-2x=12+6 by transposition 2x=18 x=18/2 by cross-multiplication x=9

OpenStudy (anonymous):

therefore the long side EF or HG is... EF=4(9)-12=36-12=24 HG=2(9)+6=18+6=24

OpenStudy (anonymous):

the shorter side FG=2x, then FG=2(9)=18 EH=FG=18

OpenStudy (anonymous):

the perimeter then is... P=EF+HG+2(FG) P=24+24+2(18) P=48+36 P=84 units